# Vision-Language Models can Identify Distracted Driver Behavior from Naturalistic Videos

### 04 Jan 2024 v3

@@ III-A Proposed Frameworks
We fundamentally propose two methodologies: (i) frame-based models (Zero-shotCLIP, Single-frameCLIP, and Multi-frameCLIP) and (ii) a video-based model (VideoCLIP). An overview of these frameworks is provided in Fig. 1. CLIP [10] is a popular vision-language model. It is proficient in most visual understanding tasks since it was trained on a large variety (around 400 million) of image-text pairs from the web.
@@
The frame-based models operate by taking inputs similar to the original CLIP model and making predictions on distracted actions. As shown in Fig. 2, the model projects the input image and text pairs into a multi-modal latent space and minimizes a contrastive loss, which enables it to learn meaningful and useful visual and textual representations. Therefore, the base model can be fine-tuned for task-specific learning. Given its ability to be instructed by natural language for a broad range of downstream tasks, such as video question answering and understanding actions from video, we develop our distracted driving analysis framework based on this model. A great advantage of using a pre-trained network over other strategies is the ability to utilize the pre-existing visual representation of the language prompt, offering the zero-shot transfer. In this work, we initially leverage CLIP’s pre-trained visual embedding for zero-shot transfer on distracted driving datasets and categorize the distracted action classes without any additional training. Following this, we fine-tune a linear classifier layer built upon the original pre-trained CLIP model. The yellow block shown in Fig. 2 represents the frozen layers. We add the green block positioned on top of the yellow layers, which represents the trainable layer. This task-based fine-tuning green block involves training a straightforward classifier layer using the pre-existing embeddings from the frozen layers for a specific task. In our case, this task is to classify distracted driving actions from visual data. We evaluate the performance of the whole framework on corresponding distracted driving datasets. As per the primary model [10], we implement six CLIP visual backbones- ViT-L/14, ViT-L/14@336px, ViT-B/16, ViT-B/32, RN101 and assess each of their performance individually.
The Zero-shotCLIP, Single-frameCLIP and Multi-frameCLIP frameworks have input $I\in\mathbb{R}^{3\times H\times W}$ where $I$, $H$, and $W$ represent the input image, image height and image width, respectively. The Zero-shotCLIP makes predictions on raw frames without requiring any training. The key difference between the other two frameworks is that Single-frameCLIP makes predictions looking at a single frame and the Multi-frameCLIP makes predictions on a stack of frames conducting a majority vote at the classification layer. For both frame-based and video-based CLIP, all the layers up to the classifier layer were kept frozen and we only finetuned the last trainable (green) layer (shown in Fig. 2).
@@
The VideoCLIP model has a similar structure to the frame-based model shown in Fig. 2, but uses a different CLIP visual encoder and model input. It utilizes a branch called S3D (separable 3D CNN) [5], a video processing backbone based on 3D-CNN, which takes a video $V\in\mathbb{R}^{S\times 3\times H\times W}$ as input and encodes the video into a feature vector $F$. Here $S$, $V$, $H$, and $W$ represent the number of frames, image height and image width, respectively. In our experiments, the number of frames $S$ is set as 30, indicating that the VideoCLIP model uses 30 frames and synthesizes them into a short video clip. The rest of the training process is similar to the frame-based CLIP. In the training process, if the cosine similarity between the video feature and its corresponding text feature is high, the distance between their feature vectors is minimized. It should be highlighted that both the Multi-frameCLIP and VideoCLIP operate on successive frames, thereby capturing the temporal context of a given distracted action. A majority vote strategy is employed at the final stage by both frameworks. The key difference between these two models is illustrated in Fig. 1, with a comprehensive discussion provided in Section IV.
Fig. 3 shows the overall sampling procedure that we follow for our experiments. The blue block on the left shows some raw frames from the original video, which has a frame rate of 30 frames per second (FPS). The yellow blocks show the sampled frames at corresponding sampling rates. Due to the CLIP network’s limitations in processing numerous frames simultaneously, we sample the frames from the video data at various sampling rates and create a 30-frame-long video clip to feed into the VideoCLIP network. From Fig. 3, it could be observed that when the sampling rate is high, such as 20 FPS, the consecutive frames exhibit little semantic differences, and the video snippet (created from the 30 frames) covers less time duration. In contrast, if the sampling rate is low, the video snippet could span a longer time duration, with greater semantic differences between two consecutive frames. We perform a series of experiments to determine the best sampling rate for the video-based framework, discussed in Section IV, and the results of the experiments are explained in Section V.
Figure 3: The raw data is shown in the blue box, which contains 300 frames of an event lasting for 10 seconds. In other words, each second covers 30 frames. We sample frames with different sampling rates from the raw data, and the sampled data is shown in the yellow box. For example, when the sampling rate is two frames per second (FPS), we sample two images from 30 raw frames covered in each second. Since the raw data lasts for 10 seconds, we will have a total of 20 sampled images in the end.
@@ IV-A Dataset
The datasets used in our study are specifically chosen for their unique driver IDs or folders for each participant. This feature is crucial for conducting our ‘leave a set of driver out’ approach while testing. We evaluate our proposed frameworks on four distinct distracted driving datasets - DMD [35], SAM-DD [4], SynDD1 [65] and StateFarm [66]. These datasets serve as standard benchmarks for the analysis of distracted driving behavior, typically providing annotated images and videos that correspond to diverse distracted driving behaviors exhibited by participants. Each dataset comes with an individual driverID (SAM-DD has separate folders for each participant) for each participant.
For our experiments, we utilize the driverIDs to separate out a specific set of drivers only for testing purposes, ensuring no overlapping between training and testing splits. Therefore, our framework does not get the chance to memorize the driver’s appearance and only focuses on distracted actions. We believe this is crucial for properly validating our algorithm’s effectiveness. This approach addresses the inaccuracies encountered in previous studies where such separation was not maintained. Moreover, to ensure consistency and reliability in our experimental approach, we executed all our experiments for all the proposed models with the same experimental settings across all the datasets we used. In addition, we assess the performance of our distracted driving models through a $k$-fold cross-validation approach executed on each driver set. Then, we calculate the average accuracy across all the $k$ folds and report it as the final accuracy of the models. Moreover, we analyze the proposed frameworks’ performances for each dataset separately and report them in the subsequent sections.
IV-B SynDD1 Dataset
The SynDD1 dataset [65] is a recent addition to distracted driving datasets, offering insights into naturalistic distracted driving behaviors. The dataset includes videos of distracted driving events recorded by cabin cameras inside the vehicle. A total of 18 participants performed various in-car driving activities, which were recorded (as video) from three distinct camera angles - the dashboard camera, the rear-view mirror camera, and the right-side window camera. The videos have two types of annotation files - gaze activity and distracted activity. The first row of Fig. 4 shows some sample frames of distracted driving actions from the dashboard camera view. There are 18 driving activities performed by the participants, including ‘normal forward driving’, ‘drinking’, ‘phone call (right hand)’, ‘phone call (left hand)’, ‘eating’, ‘texting (right hand)’, ‘texting (left hand)’, ‘hair and makeup’, ‘reaching behind’, ‘adjusting control panel’, ‘pick up something from the floor (driver seat)’, ‘pick up something from the floor (passenger seat)’, ‘talking to the passenger at right’, ‘talking to the passenger at the backseat’, ‘yawning’, ‘hand on the head’, ‘singing’, ‘shaking or dancing’. Each video file comes with annotations, including action types, the start and end time of an action, the appearance of the participant and so on. Therefore, we have the start and end time of a particular distracted action and their corresponding time duration. Each distracted driving video has a frame rate of 30 FPS. Each frame has an RGB resolution of 1920x1080 pixels. For the purpose of this study, we consider the distracted behaviors of 14 drivers, thereby generating a more concise distracted driving dataset. For further generalization, we merge the similar action categories and consider eight distinct distracted actions shown in Table I. Fig. 5 shows the distribution of the distracted driving action duration (in seconds) across eight classes. The X-axis enumerates the eight specific distracted driving action categories, while the Y-axis represents the corresponding duration (sec) of these actions aggregated across all drivers. The average action duration is approximately 18.692 seconds. The median and spread of the action duration suggest an almost uniform distribution. Consequently, the probability of a specific distracted driving action being more prevalent than others within the examined range is nearly zero. Therefore, the likelihood of occurrence for all durations of distracted driving actions is essentially equal. However, it is important to note that perfectly uniform distributions are rare to see in real-world scenarios, especially in studies related to human behavior.
Figure 4: Sample dashboard camera frames from SynDD1 dataset- (a) drinking, (b) adjusting hair, (c) talking to phone (right hand); Statefarm dataset- (d) drinking, (e) adjusting hair, (f) talking to phone (left hand)
TABLE I: Distracted driving action categories for SynDD1 dataset
ID distracted driving prompt description
0 driver is adjusting his or her hair while driving a car
1 driver is drinking water from a bottle while driving a car
2 driver is eating while driving a car
3 driver is picking something from the floor while driving a car
4 driver is reaching behind to the backseat while driving a car
5 driver is singing a song with music and smiling while driving
6 driver is talking to the phone on hand while driving a car
7 driver is yawning while driving a car
Figure 5: The distribution of the time duration (sec) across eight distracted driving actions from the SynDD1 dataset. The mean action duration is 18.692 seconds. The median and spread of the action duration suggest an almost uniform distribution, demonstrating a balance in the action categories we use.
IV-C StateFarm Dataset
The StateFarm dataset [66] serves as a common dataset for distracted driving behavior analysis, providing only image (RGB) data. The images were captured from an in-cabin camera positioned on the right-side window, offering a right-side perspective of the driver, as (d), (e) and (f) of Fig. 4. Data were collected from 26 participants, and the total number of image frames is 22,424. Similar to the SynDD1 dataset, the StateFarm dataset includes unique driver IDs for each participant, which enables us to perform the ‘leave a set of driver out’ step and driver-level k-fold cross-validation. However, it does not provide information on the action duration, start, or end times of actions as the labels pertain only to image frames. There are ten distracted driving activity categories in total- ‘normal driving’, ‘texting (right hand)’, ‘talking on the phone (right hand)’, ‘texting (left hand)’, ‘talking on the phone (left hand)’, ‘operating the radio’, ‘drinking’, ‘reaching behind’, ‘hair and makeup’, ‘talking to a passenger’. The images have a resolution of (640,480,3). We consider all the drivers and ten action categories for our experimental analysis.
IV-D DMD Dataset
The Driver Monitoring Dataset (DMD) is a specialized dataset  [35] designed to facilitate research in driver distraction and behavior recognition. This dataset captures various aspects of driver activity in 37 participants- focusing on distraction, drowsiness, hand movements, and gaze direction. This dataset offers RGB images (up to 1920x1080 pixels), depth data, and corresponding metadata with ten distraction categories. The data was recorded in two different environments - outdoors in a real car and indoors in a driving simulator. There are ten distracted driving actions, including ‘drinking water’, ‘adjusting hair and makeup’, ‘talking to the phone on right hand’, ‘adjusting radio’, ‘reaching to the backseat’, ‘reaching side’, ‘driving safely’, ‘talking to the passenger’, ‘texting on the phone with right hand’ and ‘yawning’. We implemented a 7-fold cross-validation on the ten distraction classes with 78917 samples for each fold and 552419 samples in total for all the folds. In addition, we performed the ‘leave a set of drivers out’ method using the provided driver IDs within the 7-fold cross-validation to evaluate our model performance.
IV-E SAM-DD Dataset
The Singapore AutoManNTU distracted driving (SAM-DD) dataset [4] is a comprehensive dataset for intelligent driving research. It includes image data from 42 participants (34 males, 8 females) of diverse ages and driving experiences. It features 51,175 samples across front and side camera views, encompassing both RGB and Depth image modes. The dataset includes ten distinct distraction classes with nine physically distracted behaviors - ‘drinking’, ‘talking’, ‘texting’, ‘touching hair’, ‘adjusting glasses’, ‘reaching behind’, and a fatigue-related behavior (head dropping). The high resolution (1200x900 pixels) RGB data makes it suitable for developing learning-based models in various driving environments and driver state-related applications. Both the DMD and SAM-DD are large datasets in terms of distracted class, number of samples per class and number of drivers. To ensure consistency and reliability in our experimental approach, we executed our proposed models with the same experimental settings for all the datasets.
IV-F Experiment Setup - ZeroshotCLIP and Single-frameCLIP
Our overall approach is depicted in Fig. 2. First, we explore the zero-shot transfer capability of the CLIP model on distracted driving data in the Zero-shotCLIP framework. Then, we explore the task-based finetuning route to evaluate the “task learning capability” of our proposed Single-frameCLIP framework. For the Zero-shotCLIP experiment, there is no training associated, and we simply predict the class-level confidence score of the pre-trained CLIP model on raw image frames (224x224x3). Next, we assess the representation learning capability of the pre-trained CLIP model on the distracted driving datasets in the Single-frameCLIP framework. We add an extra linear classifier layer on top of the pre-trained CLIP model architecture and train the single layer on distracted driving datasets. For training the linear classifier (see CLIP [10]), we use image frames of input size (224x224x3) collected from the four datasets. We utilize a batch size of 100 and a learning rate of 0.001. While training the Single-frameCLIP, we kept the pre-trained layers frozen and updated only the final layer based on the cross-entropy loss. Fig. 1 provides a visual representation of this approach, which generates predictions on a frame-by-frame basis. To illustrate this process, let’s consider a 10-second event sampled at a rate of 30 frames per second with a total of 300 frames (depicted in Fig. 1 by the yellow box). For each of these frames, the Single-frameCLIP model generates a prediction, beginning with the first frame (Image-1) and subsequently for each following frame (Image-2, Image-3, etc). In addition, based on the driverID, we perform an 8-fold cross-validation for the StateFarm datasets and a 7-fold cross-validation for the other three datasets. We report the average accuracy in Section V. Moreover, we experiment with various CLIP backbones (see CLIP [10]) and other hyperparameters. Notably, the original CLIP model employs semantic information in text labels rather than the traditional one-hot labels, which require providing a textual description or prompt for each action category. Therefore, we provide a full sentence as a prompt for describing the image frame of each action class instead of a single word. The prompt descriptions for the SynDD1 dataset are shown in Table I. As reported in [10], we consider manually fine-tuning and rephrasing the prompt sentences for our experiments, an area that can be further optimized in future work.
IV-G Experiment Setup - Multi-frameCLIP
The Multi-frameCLIP framework is specifically designed to analyze sequences of consecutive frames from distracted driving videos. This approach extends upon the Single-frameCLIP, concluding with an additional majority voting step applied across consecutive frames to generate the final event prediction. The majority voting method allows our framework to make a single prediction on a particular action event. During the first step of training, this framework is provided with the frames between the start and end times of a specific action within the video footage. Following this, the Single-frameCLIP model is applied to each frame falling within these timestamps, providing individual predictions for each frame. A majority voting method is then employed to produce a final prediction. If this prediction aligns with the ground truth action label for that segment of the video, it is considered a correct prediction. This entire process is visually represented in Fig. 1. From the raw video data, we sample frames at varying rates (as depicted in Fig. 3). Let’s consider a 10-second event containing 300 frames captured at a rate of 30 FPS (shown by the yellow box). The Single-frameCLIP model is employed to make predictions for each of these frames, starting with the first frame (Image-1) and then sequentially progressing (Image-2, Image-3, and so on). Finally, a majority vote among all the individual frame predictions is taken to make a final prediction for the entire 10-second event in the video. This method allows Multi-frameCLIP to extend Single-frameCLIP’s capability by incorporating a temporal understanding of the events in a video clip. To ensure robust results, we conducted cross-validation experiments with sampling rates of 1, 2, 5, 20, and 0.5 frames per second.
IV-H Experiment Setup - VideoCLIP
The VideoCLIP is an adaptation of the frame-based CLIP model. It incorporates a 3D Convolutional Neural Network (CNN) backbone specifically designed for video processing. For additional information, readers can refer to [67]. The key difference in this model is its consideration of the temporal information of video data when making a prediction. When the model needs to make a prediction about a particular frame, it does not just look at that single frame. Instead, it combines 14 frames that come before and 15 frames that come after a particular target frame to form a small video footage. A total of 30 frames (including the target frame) are used together for prediction. This video piece is fed into the pre-trained CLIP backbone and a majority voting step is conducted to produce the final prediction. Fig. 1 shows an overview of the VideoCLIP model pipeline. This technique enables the model to consider not only the content of the targeted frame but also the preceding and subsequent frames. In this way, the network can perform classification with more complete temporal information. In this experiment, the learning rate and batch size are set as 0.0003 and 5, respectively, and Adam [68] optimizer is used. We followed the ‘leave a set of driver out’ method as before and split the drivers into two non-overlapping sets (training and testing) to evaluate the performance of never-before-seen drivers. Finally, we perform a 7-fold cross-validation on the splits with sampling rates of 1,2,5 and 20 frames per second.
IV-I Experiment Setup for CLIP vs. Other Models
For performance comparison, we further trained several mainstream CNN backbones, which are commonly used for distracted driving recognition. For this experiment, we considered five different backbones - AlexNet [43], VGG16 [41], MobileNet-v2 [45], ResNet18 and ResNet50 [42], and implemented them on the frames sampled from the video data. We used the ImageNet pre-trained weights and fine-tuned the models with a learning rate of 0.001 and a batch size of 64. We used Adam [68] as the optimizer. We report the experimental results and outcome of our observation in Section V.

### 21 Mar 2024 v4

@@ III-A Proposed Frameworks
We fundamentally propose two types of models: (i) frame-based models (Zero-shotCLIP, Single-frameCLIP, and Multi-frameCLIP) and (ii) video-based models (VideoCLIP). An overview of these frameworks is provided in Fig. 1. CLIP [11] is a popular vision-language model. It is proficient in most visual understanding tasks since it was trained on a large variety (around 400 million) of image-text pairs from the internet.
@@
The frame-based models operate by taking inputs similar to the original CLIP model and making predictions on distracted actions. As shown in Fig. 2, the model projects the input image and text pairs into a multi-modal latent space and minimizes a contrastive loss, which enables it to learn meaningful and useful visual and textual representations. Therefore, the base model can be fine-tuned for task-specific learning. Given its ability to be instructed by natural language for a broad range of downstream tasks, such as video question answering and understanding actions from video, we develop our distracted driving analysis framework based on this model. A great advantage of using a pre-trained network over other strategies is the ability to utilize the pre-existing visual representation of the language prompt, offering the zero-shot transfer. In this work, we initially leverage CLIP’s pre-trained visual embedding for zero-shot transfer on distracted driving datasets and categorize the distracted action classes without any additional training. Following this, we fine-tune a linear classifier layer built upon the original pre-trained CLIP model. The yellow block shown in Fig. 2 represents the frozen layers. We add the green block positioned on top of the yellow layers, which represents the trainable layer. This task-based fine-tuning green block involves training a straightforward classifier layer using the pre-existing embeddings from the frozen layers for a specific task. In our case, this task is to classify distracted driving actions from visual data. We evaluate the performance of the whole framework on corresponding distracted driving datasets. As per the primary model [11], we implement five CLIP visual backbones- ViT-L/14, ViT-L/14@336px, ViT-B/16, ViT-B/32, RN101 and assess their performance individually.
The Zero-shotCLIP, Single-frameCLIP and Multi-frameCLIP frameworks have input $I\in\mathbb{R}^{3\times H\times W}$ where $I$, $H$, and $W$ represent the input image, image height and image width, respectively. The Zero-shotCLIP makes predictions on raw frames without requiring any training. The key difference between the other two frameworks is that Single-frameCLIP makes predictions looking at a single frame and the Multi-frameCLIP makes predictions on a stack of frames conducting a majority vote at the classification layer. For both frame-based and video-based CLIP, all the layers up to the classifier layer are kept frozen and we only finetune the last trainable layer (the green layer in Fig. 2).
@@
The VideoCLIP model has a similar structure to the frame-based model shown in Fig. 2, but uses a different CLIP visual encoder and model input. It utilizes a branch called S3D (separable 3D CNN) [75], a video processing backbone based on 3D-CNN, which takes a video $V\in\mathbb{R}^{S\times 3\times H\times W}$ as input and encodes the video into a feature vector $F$. Here $S$, $V$, $H$, and $W$ represent the number of frames, image height and image width, respectively. In our experiments, the number of frames $S$ is set as 30, indicating that the VideoCLIP model uses 30 frames and synthesizes them into a short video clip. The rest of the training process is similar to the frame-based CLIP. In the training process, if the cosine similarity between the video feature and its corresponding text feature is high, the distance between their feature vectors is minimized. It should be highlighted that both the Multi-frameCLIP and VideoCLIP operate on successive frames, thereby capturing the temporal context of a given distracted action. A majority vote strategy is employed at the final stage by both frameworks. The key difference between these two models is illustrated in Fig. 1.
@@ IV-A Dataset
In this study, we utilize four common distracted driving datasets—DMD [12], SAM-DD [18], SynDD1 [76] and StateFarm [77] to investigate driver behavior under various real-world conditions. These datasets (summarized in Table I) serve as the foundation for evaluating our proposed frameworks, offering a wide variety of annotated images and videos with diverse distracted driving scenarios. The action categories span common distractions that affect driving safety, providing a comprehensive resource for analyzing distracted driving behavior. In addition, they incorporate a range of modalities including RGB, IR, and Depth, with input resolutions up to 1920x1080. The datasets used in our study are specifically chosen for their unique driver IDs for each participant. It is crucial to conduct the ‘leave a set of driver out’ evaluation.
The usage of these diverse datasets ensures the generalizability of our proposed framework by incorporating a wide range of distraction actions, data modalities, input resolutions, and various camera views to capture comprehensive driving behaviors across different scenarios and conditions. A significant number of drivers (ranging from 18 to 42) with varied driving habits and physical characteristics further enriches the data pool, enabling our framework to learn from a broad spectrum of real-world driving situations. This diversity allows us to develop robust and adaptable frameworks capable of accurately identifying distracted driving behaviors across different vehicles, environments, and driver demographics. As a result, our frameworks can be optimized to perform effectively in real-world applications, reducing the risk of overfitting to specific conditions and enhancing their applicability across a wide range of driving contexts.
TABLE I: Overview of Distracted Driving Datasets
Dataset Modality Input Data No. of No. of No. of Action
name resolution type camera distraction drivers categories
view actions
DMD [12]
RGB
IR
Depth
1920x1080
Image
Video
2 10 37
drinking, hair and makeup, talking on phone (right)
operating radio, reaching behind, reaching side, safe driving
talking to passenger, texting (right), yawning
SAM-DD [18] RGB 1200x900
Image
Video
2 10 42
safe driving, drinking, talking on phone (left),
talking on phone (right), texting (left), texting (right),
hair, adjusting glasses, reaching behind, head dropping
StateFarm [77] RGB 640x480 Image 1 10 26
safe driving, texting (right), talking on phone (right),
texting (left), talking on phone (left), operating radio, drinking,
reaching behind, hair and makeup, talking to passenger
SynDD1 [76] RGB 1920x1080
Image
Video
3 8 18
adjusting hair, drinking, eating,
picking up something, reaching behind,
singing, talking on phone, yawning
TABLE II: Summary of Experimental setup
Framework Approach Input Training Cross Leaving a set
name resolution details validation of driver out
Zero-shotCLIP
Zero-shot
transfer
224x224x3
No training,
use pre-trained CLIP
N/A N/A
Single-frameCLIP
Fine-tuning
classifier-layer
224x224x3
Training linear classifier layer
batch size: 100, optimizer: LBFGS
8-fold (StateFarm),
7-fold (others)
Multi-frameCLIP
Sequence-analysis
with majority voting
variable
Single-frameCLIP predictions
aggregated with majority voting
7-fold
VideoCLIP
Video feature-analysis
with temporal context
224x224x3x30
(30 frames)
Learning rate: 0.0003
batch size: 5, optimizer: Adam
7-fold
IV-B Experiment Setup
Table II summarizes our experimental setup across the image-based and video-based models. It shows how we leverage the versatility of the CLIP model to learn distracted driving patterns, evolving from zero-shot learning to advanced video analysis. First, we combine the pre-trained CLIP backbones with sentence-level prompts in Zero-shotCLIP to get predictions on raw images without prior training. Then, the Single-frameCLIP extends this idea by fine-tuning an additional classifier layer on our distracted driving datasets. Finally, Multi-frameCLIP applies a majority voting on top of Single-frameCLIP logits and VideoCLIP incorporates a 3D CNN backbone [75] on consecutive frames to make a “distraction event” prediction. We maintain the standard CLIP input resolution shown in the second column of the table.
The original CLIP model fine-tuning requires sentence descriptions as labels rather than the traditional single text label. Therefore, we provide a full sentence description as a prompt for each action category. As reported in [11], we consider manually fine-tuning and rephrasing the prompts for our experiments, an area that can be further optimized in future research. While testing various hyperparameter settings (backbones, FPS, camera-view etc.), we select the parameters that deliver the highest performance for our proposed models and refer to them as the “optimal hyperparameter setting”. Furthermore, to compare the performance of our frameworks with traditional baselines, we trained five CNN backbones (AlexNet [50], VGG16 [48], MobileNet-v2 [52], ResNet18 and ResNet50 [49]) on the same dataset. For this experiment, we load the ImageNet pre-trained weights to train these models with a learning rate of 0.001, a batch size of 64 and Adam [78] as the optimizer. Moreover, to resolve the issue raised in [33], we design the experiments to leave a specific subset of drivers separate only for testing rounds, ensuring a strict separation between training and testing splits. In addition, we adopt a $k$-fold cross-validation method on each driver set, where the final model accuracy is assessed by averaging accuracies over all folds. Furthermore, to maintain consistency and reliability in our experimental assessments, we maintain uniform experimental settings across all models and datasets. The results of these experiments along with the effect of the optimal hyperparameter setting are presented in Section V.